# Group-weighted conformal prediction

### v2

@@ 1 Introduction
The conformal prediction (CP) framework offers a method for constructing prediction intervals, for the problem of predicting a response variable $Y\in\mathcal{Y}$ based on features $X\in\mathcal{X}$, in a distribution-free way (see Vovk et al., (2005) for background). In its standard form, conformal prediction deals with the case of i.i.d. data (or, more generally, exchangeable data)—namely, the training data used for constructing the prediction intervals, and the test data on which the prediction intervals are deployed, are assumed to be drawn i.i.d. from the same distribution.
To relax this assumption, Tibshirani et al., (2019) extend conformal prediction to the setting of covariate shift, where the test data distribution $Q$ may differ from the training data distribution $P$ in terms of the marginal distribution of the covariates—that is, $Q_{X}$ may differ from $P_{X}$ (but the dependence of the response $Y$ given features $X$ is assumed to be the same across training and test data, i.e., the conditional distributions $P_{Y|X}$ and $Q_{Y|X}$ are assumed to be equal). However, the resulting method, weighted conformal prediction (WCP), requires knowledge of the covariate shift (specifically, the ratio between the marginal distributions $P_{X}$ and $Q_{X}$, used to reweight the training data to better resemble the test data distribution). For example, if $P_{X}$ and $Q_{X}$ have densities $f$ and $g$, respectively, implementing WCP requires access to a weight function $w(x)\propto g(x)/f(x)$. However, in a real data setting, it is not realistic to assume that $w(x)$ is known exactly. In practice the weight function would instead be approximated by some estimate $\widehat{w}(x)$, e.g., by estimating these unknown densities via fitting models to the training and test data. The error in this estimation process—i.e., the error of $\widehat{w}$ in estimating $w$—can then lead to a failure of coverage in the resulting prediction intervals.
@@ 2.1 Conformal prediction
The conformal prediction (CP)framework, which aims to provide prediction intervals with distribution-free validity, was initially developed beginning in the late 1990’s (see, e.g., early works by Gammerman et al., (1998); Saunders et al., (1999); Papadopoulos et al., (2002), and see Vovk et al., (2005) for a comprehensive overview). The CP methodology quickly gained popularity within the statistics community as a robust tool for regression and classification (Lei et al.,, 2018; Romano et al.,, 2019). Conformal prediction is a procedure that can be used in conjunction with any regression procedure or modeling algorithm, to produce prediction intervals with valid marginal coverage guarantees.
In this section and throughout the paper, we will restrict our attention to one particular version of this procedure, the split conformal method (also called inductive conformal) (Vovk et al.,, 2005; Papadopoulos,, 2008), which separates the process into two steps, model fitting and calibration, each run on a different portion of the data set—this offers a computationally efficient option (relative to the high computational cost of the alternative, full conformal, which we do not cover here). For convenience, we assume that the training data consists of $n^{*}+n$ many data points, partitioned into data sets $\{(X^{*}_{i},Y^{*}_{i})\}_{i\in[n^{*}]}$ (used for model fitting) and $\{(X_{i},Y_{i})\}_{i\in[n]}$ (used for calibration). We will refer to these two data sets as the pretraining set and the calibration set. (Of course, in practice we would be given a single data set and would then partition it randomly into two parts, i.e., data splitting, but the notation of two separate data sets will be more convenient for the exposition of the paper.) The split conformal procedure then operates as follows:
@@ 2.2 Weighted conformal prediction and covariate shift
To address this setting, Tibshirani et al., (2019) proposed a modification of CP, the weighted conformal prediction (WCP) method. Working again in the split conformal framework (so that we can assume a score function $s=s(x,y)$ has been pretrained), the WCP method modifies the calibration step by computing a weighted quantile of scores:
@@
where $\delta_{s}$ is the point mass at $s$, and where $w(x)$ captures the shift from the training feature distribution $P_{X}$ to the test feature distribution $Q_{X}$. The prediction interval is then again given by $\widehat{C}_{n}(x)=\left\{y\in\mathcal{Y}:s(x,y)\leq\widehat{q}\right\}$, as for the unweighted case. (To compare to the unweighted version of the method, we can observe that the quantile $\widehat{q}$ computed in (1) can equivalently be defined as
$\widehat{q}=\textnormal{Quantile}_{1-\alpha}\left(\sum_{i=1}^{n}\frac{1}{n+1}\cdot\delta_{s_{i}}+\frac{1}{n+1}\cdot\delta_{+\infty}\right),$
which coincides with the definition of $\widehat{q}$ for the WCP method if we define a constant weight function $w(x)\equiv 1$—that is, in the case $P_{X}=Q_{X}$, where there is no covariate shift.)
@@
The WCP method has subsequently been applied to a range of statistical problems that can be reformulated as instances of covariate shift, including applications to survival analysis (Gui et al.,, 2022; Candès et al.,, 2023), causal inference (Lei and Candès,, 2021; Yin et al.,, 2022; Jin et al.,, 2023), and adaptive learning (Fannjiang et al.,, 2022).
@@
In practice, a major challenge in the implementation of the WCP procedure is that the weight function $w(x)\propto\frac{\mathsf{d}Q_{X}}{\mathsf{d}P_{X}}(x)$ is not known exactly—indeed, Tibshirani et al., (2019)’s initial results for WCP offer no theoretical guarantees for an estimated $w(x)$, although their empirical results demonstrate that estimating $w$ on a large sample performs well in practice. More recent results in the literature offer guarantees that quantify the extent to which errors in estimating $w$ can lead to loss of coverage. For example, Lei and Candès, (2021) prove the following guarantee. Assume we are given a fixed function $\widehat{w}(x)$ (or more generally, a function $\widehat{w}(x)$ that is estimated independently of the calibration data and test point and can thus be treated as fixed—for instance, $\widehat{w}(x)$ may be fitted on the training set). If the true and estimated weights are normalized to satisfy $\mathbb{E}_{P_{X}}[w(X)]=\mathbb{E}_{P_{X}}[\widehat{w}(X)]=1$, then the WCP predictive interval satisfies
$\mathbb{P}_{P^{n}\times Q}\left\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\right\}\geq 1-\alpha-\mathbb{E}\left[\frac{\left|\widehat{w}(X)-w(X)\right|}{2}\right].$ (3)
Related results, also establishing bounds on the loss of coverage in terms of the difference $\widehat{w}-w$, can be found throughout the literature, e.g., see Candès et al., (2023); Gui et al., (2022); Yang et al., (2022); Jin et al., (2023); Yin et al., (2022).
@@ 3 Problem setting: group-wise covariate shift
if $X^{0}=k$, for each $k\in[K]$. In particular, implementing WCP (and thus ensuring a predictive coverage guarantee) relies only on knowing these weights; coverage does not rely on any knowledge or assumptions for the distributions $\Pi_{k}$, which determine the feature and response distributions within each group $k$, since these are assumed to be the same across the training and test data.
Throughout this paper, we will assume that the test proportions $q_{k}$ are known. Effectively, we are assuming that there is ample unlabeled data (i.e., test points, for which we observe $X$ but not $Y$), so that we can estimate the proportion of subgroups within the target population with extremely high accuracy. Alternatively we can interpret this assumption in a different way: we are simply requiring that coverage holds relative to some prespecified distribution over the $K$ groups (e.g., if we set $q_{k}\equiv 1/K$, this indicates that we are interested in the average group-wise coverage). The problem of estimating a weight function $w(X)$ then reduces to the question of estimating the proportions $p_{1},\dots,p_{K}$, to characterize the sampling bias of the training data, drawn from $P$, as compared to the target distribution $Q$.
@@ 3.1 Group-weighted conformal prediction (GWCP)
In other words, the prediction interval is defined via a score threshold given by the $(1-\alpha)$-quantile of the following distribution: compute the empirical distribution $\widehat{P}_{\textnormal{score}}^{(k)}$ of scores from data points in group $k$, and then take the mixture of these empirical distributions by placing weight $q_{k}$ on the $k$th component, for each group $k\in[K]$.
@@
We will now see that we can view this as a slightly less conservative version of the WCP method. To see why, suppose we let $\widehat{p}_{k}=n_{k}/n$—that is, if the calibration data points $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ are sampled from the $K$ groups with probabilities $p_{1},\dots,p_{K}$, as in (4), then $\widehat{p}_{k}$ is an estimate of $p_{k}$ for each $k$. Then, for any $X$ with $X^{0}=k$, the weight for the covariate shift under this model is given by $w(X)=q_{k}/p_{k}$; we can estimate this weight with $\widehat{w}(X)=q_{k}/\widehat{p}_{k}$. (In this comparison, we will assume for simplicity that $n_{k}>0$ for all $k$, i.e., all groups are observed in the training set—otherwise the function $\widehat{w}$ would not be well-defined.) By definition, we have $\sum_{i\in[n]}\widehat{w}(X_{i})=n$. Then rewriting the definition of $\widehat{q}$ one more time,
@@
As we will see below, the prediction interval defined here in (6) leads to a coverage guarantee that is slightly smaller than the target level $1-\alpha$. Interestingly, however, this cannot be fixed by returning to the slightly more conservative WCP method given in (2). We choose to work with this present formulation, (6), since avoiding the additional weight on $\delta_{+\infty}$ leads to a simpler construction and a (slightly) less conservative method without any additional loss of coverage in the theoretical guarantee. (Of course, any coverage guarantee that can be established for the prediction interval defined in (6) will also hold for the version of WCP that adds a weight to $\delta_{+\infty}$, since this leads to a strictly more conservative prediction interval.)
@@ 3.2 Estimation error in the group-wise setting: applying prior results
However, even in this setting of a finite-dimensional unknown parameter, existing coverage guarantees provide a fairly pessimistic view and imply that the loss of coverage may be severe. In particular, if our estimate of the probabilities $p_{1},\dots,p_{K}$ (which determine the relative proportions of the $K$ groups in the training distribution $P$) is based on a sample of size $\mathcal{O}(n)$, then we would expect that each $p_{k}$ is estimated with error scaling as $\mathcal{O}(\sqrt{p_{k}/n})$ (due to the variance of a Binomial distribution); equivalently, the relative error in estimating $p_{k}$ scales as $\mathcal{O}(1/\sqrt{np_{k}})$. For example, if $p_{k}\equiv 1/K$ (i.e., the true distribution places equal weight on each group), then the relative error in estimating each $p_{k}$ scales as $\mathcal{O}(\sqrt{K/n})$, and so the coverage loss term in Lei and Candès, (2021)’s guarantee (3) can be expected to scale as
@@
To demonstrate this intuition more concretely, we compute Lei and Candès, (2021)’s coverage guarantee (3) empirically, across a range of sample sizes $n=100,110,120,\dots,1000$, in three regimes: a constant number of groups, $K=10$; a slowly increasing number of groups, $K=\lfloor\sqrt{n}\rfloor$; and a proportional number of groups, $K=n/10$. In each case, the probabilities are given by $p_{k}=q_{k}\equiv 1/K$ for all groups $k\in[K]$, for both the training and test distributions. The estimated weight function $\widehat{w}$ is then given by $\widehat{w}(X)=q_{k}/\widehat{p}_{k}$ for any data point with $X^{0}=k$, where $q_{k}\equiv 1/K$ is known, while $\widehat{p}_{k}$ is given by the empirical proportion of group $k$ in a pretraining sample of size $n$. To avoid dividing by zero, we add a +1 adjustment to each group’s count—that is, $\widehat{p}_{k}=\frac{n_{k}+1}{n+K}$, where $n_{k}$ is the number of samples observed in group $k$ within the training set of size $n$ (since the true proportions $p_{k}$ are uniform, this adjustment can only make the estimates more accurate).
@@
Finally, we mention another line of the existing literature, which seeks to provide group-conditional coverage type guarantees (i.e., in the notation of this work, guarantees of the type $\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\mid X_{n+1}^{0}=k\}$). For instance, works such as Dunn et al., (2022); Ding et al., (2023) study problems related to this framework. Any guarantee of group-conditional coverage will automatically yield coverage with respect to a group-wise covariate shift—that is, this type of guarantee is strictly stronger than coverage with respect to Q, which we study in this paper. However, a group-conditional guarantee is meaningfully achievable only when within-group sample size, $n_{k}$, is fairly large, while our work allows for small $n_{k}$ (as long as $K$, the number of groups, is large).
Figure 1: Lei and Candès, (2021)’s WCP coverage guarantee (3) for varying numbers of groups. The horizontal line marks the target coverage level 90% (i.e., $\alpha=0.1$). See Section 3.2 for details.

### v4

@@ 1 Introduction
The conformal prediction (CP) framework offers a method for constructing prediction intervals, for the problem of predicting a response variable $Y\in\mathcal{Y}$ based on features $X\in\mathcal{X}$, in a distribution-free way (see (Vovk et al., 2005) for background). In its standard form, conformal prediction deals with the case of i.i.d. data (or, more generally, exchangeable data)—namely, the training data used for constructing the prediction intervals, and the test data on which the prediction intervals are deployed, are assumed to be drawn i.i.d. from the same distribution.
To relax this assumption, (Tibshirani et al., 2019) extend conformal prediction to the setting of covariate shift, where the test data distribution $Q$ may differ from the training data distribution $P$ in terms of the marginal distribution of the covariates—that is, $Q_{X}$ may differ from $P_{X}$ (but the dependence of the response $Y$ given features $X$ is assumed to be the same across training and test data, i.e., the conditional distributions $P_{Y|X}$ and $Q_{Y|X}$ are assumed to be equal). However, the resulting method, weighted conformal prediction (WCP), requires knowledge of the covariate shift (specifically, the ratio between the marginal distributions $P_{X}$ and $Q_{X}$, used to reweight the training data to better resemble the test data distribution). For example, if $P_{X}$ and $Q_{X}$ have densities $f$ and $g$, respectively, implementing WCP requires access to a weight function $w(x)\propto g(x)/f(x)$. However, in a real data setting, it is not realistic to assume that $w(x)$ is known exactly. In practice the weight function would instead be approximated by some estimate $\widehat{w}(x)$, e.g., by estimating these unknown densities via fitting models to the training and test data. The error in this estimation process—i.e., the error of $\widehat{w}$ in estimating $w$—can then lead to a failure of coverage in the resulting prediction intervals.
@@ 2.1 Conformal prediction
The conformal prediction (CP)framework, which aims to provide prediction intervals with distribution-free validity, was initially developed beginning in the late 1990’s (see, e.g., early works by (Gammerman et al., 1998; Saunders et al., 1999; Papadopoulos et al., 2002), and see (Vovk et al., 2005) for a comprehensive overview). The CP methodology quickly gained popularity within the statistics community as a robust tool for regression and classification (Lei et al., 2018; Romano et al., 2019). Conformal prediction is a procedure that can be used in conjunction with any regression procedure or modeling algorithm, to produce prediction intervals with valid marginal coverage guarantees.
In this section and throughout the paper, we will restrict our attention to one particular version of this procedure, the split conformal method (also called inductive conformal) (Vovk et al., 2005; Papadopoulos, 2008), which separates the process into two steps, model fitting and calibration, each run on a different portion of the data set—this offers a computationally efficient option (relative to the high computational cost of the alternative, full conformal, which we do not cover here). For convenience, we assume that the training data consists of $n^{*}+n$ many data points, partitioned into data sets $\{(X^{*}_{i},Y^{*}_{i})\}_{i\in[n^{*}]}$ (used for model fitting) and $\{(X_{i},Y_{i})\}_{i\in[n]}$ (used for calibration). We will refer to these two data sets as the pretraining set and the calibration set. Of course, in practice we would be given a single data set and would then partition it randomly into two parts, i.e., data splitting, but the notation of two separate data sets will be more convenient for the exposition of the paper. The split conformal procedure then operates as follows:
@@ 2.2 Weighted conformal prediction and covariate shift
To address this setting, (Tibshirani et al., 2019) proposed a modification of CP, the weighted conformal prediction (WCP) method. Working again in the split conformal framework (so that we can assume a score function $s=s(x,y)$ has been pretrained), the WCP method modifies the calibration step by computing a weighted quantile of scores:
@@
where $\delta_{s}$ is the point mass at $s$, and where $w(x)$ captures the shift from the training feature distribution $P_{X}$ to the test feature distribution $Q_{X}$. The prediction interval is then again given by $\widehat{C}_{n}(x)=\left\{y\in\mathcal{Y}:s(x,y)\leq\widehat{q}\right\}$, as for the unweighted case. To compare to the unweighted version of the method, we can observe that the quantile $\widehat{q}$ computed in (1) is a special case of  (2) with a constant weight function $w(x)\equiv 1$—that is, in the case $P_{X}=Q_{X}$, where there is no covariate shift.
@@
The WCP method has subsequently been applied to a range of statistical problems that can be reformulated as instances of covariate shift, including applications to survival analysis (Gui et al., 2022; Candès et al., 2023), causal inference (Lei and Candès, 2021; Yin et al., 2022; Jin et al., 2023), and adaptive learning (Fannjiang et al., 2022).
@@
In practice, a major challenge in the implementation of the WCP procedure is that the weight function $w(x)\propto\frac{\mathsf{d}Q_{X}}{\mathsf{d}P_{X}}(x)$ is not known exactly—indeed, (Tibshirani et al., 2019)’s initial results for WCP offer no theoretical guarantees for an estimated $w(x)$, although their empirical results demonstrate that estimating $w$ on a large sample performs well in practice. More recent results in the literature offer guarantees that quantify the extent to which errors in estimating $w$ can lead to loss of coverage. For example, (Lei and Candès, 2021) prove the following guarantee. Assume we are given a fixed function $\widehat{w}(x)$ (or more generally, a function $\widehat{w}(x)$ that is estimated independently of the calibration data and test point and can thus be treated as fixed—for instance, $\widehat{w}(x)$ may be fitted on the training set). If the true and estimated weights are normalized to satisfy $\mathbb{E}_{P_{X}}[w(X)]=\mathbb{E}_{P_{X}}[\widehat{w}(X)]=1$, then the WCP predictive interval satisfies
$\mathbb{P}_{P^{n}\times Q}\left\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\right\}\geq 1-\alpha-\mathbb{E}_{P_{X}}\left[\frac{\left|\widehat{w}(X)-w(X)\right|}{2}\right].$ (3)
Related results, also establishing bounds on the loss of coverage in terms of the difference $\widehat{w}-w$, can be found throughout the literature, e.g., see Candès et al. (2023); Gui et al. (2022); Yang et al. (2022); Jin et al. (2023); Yin et al. (2022).
An alternative approach: within-group coverage.
Finally, we mention another line of the existing literature, which seeks to provide group-conditional coverage type guarantees (i.e., in the notation of this work, guarantees of the type $\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\mid X_{n+1}^{0}=k\}$). For instance, works such as (Dunn et al., 2022; Ding et al., 2023) study problems related to this framework. Any guarantee of group-conditional coverage will automatically yield coverage with respect to a group-wise covariate shift—that is, this type of guarantee is strictly stronger than coverage with respect to Q, which we study in this paper. However, a group-conditional guarantee is meaningfully achievable only when within-group sample size, $n_{k}$, is fairly large, while our work allows for small $n_{k}$ (as long as $K$, the number of groups, is large). Additional related approaches in the literature include the work of (Gupta et al., 2020) on bin-wise calibration, (Gibbs et al., 2023),(Jung et al., 2022) which focus on group conditional coverage guarantees.
@@ 3 Problem setting: group-wise covariate shift
if $X^{0}=k$, for each $k\in[K]$. In particular, implementing WCP and thus ensuring a predictive coverage guarantee relies only on knowing these weights; coverage does not rely on any knowledge or assumptions for the distributions $\Pi_{k}$, which determine the feature and response distributions within each group $k$, since these are assumed to be the same across the training and test data.
For most of this paper, we will assume that the test proportions $q_{k}$ are known. Effectively, we are assuming that there is ample unlabeled data i.e., test points, for which we observe $X$ but not $Y$, so that we can estimate the proportion of subgroups within the target population with extremely high accuracy or we simply want coverage with respect to the empirical assignment of test points to the different groups. In Section 5.3, we will extend to the setting where these proportions are unknown, and the $q_{k}$’s are instead estimated empirically. Alternatively we can interpret the assumption that the $q_{k}$’s are known in a different way: we are simply requiring that coverage holds relative to some prespecified distribution over the $K$ groups. For example, if we set $q_{k}\equiv 1/K$, this indicates that we are interested in the average group-wise coverage. The problem of estimating a weight function $w(X)$ then reduces to the question of estimating the proportions $p_{1},\dots,p_{K}$, to characterize the sampling bias of the training data, drawn from $P$, as compared to the target distribution $Q$.
@@ 3.1 Group-weighted conformal prediction (GWCP)
In other words, the prediction interval is defined via a score threshold given by the $(1-\alpha)$-quantile of the following distribution: compute the empirical distribution $\widehat{P}_{\textnormal{score}}^{(k)}$ of scores from data points in group $k$, and then take the mixture of these empirical distributions by placing weight $q_{k}$ on the $k$th component, for each group $k\in[K]$. A useful feature of this construction is that we do not require the group assignments of the points in the test set to be known.
Dunn et al. (2022) used this kind of CDF pooling technique in the context of a two-layer hierarchical model, but only asymptotic guarantees are shown; this type of approach is also studied by Lee et al. (2023), who show finite sample guarantees, but again in a different setting where new groups are sampled from a hierarchical model.
@@
We will now see that we can view this as a slightly less conservative version of the WCP method. To see why, suppose we let $\widehat{p}_{k}=n_{k}/n$—that is, if the calibration data points $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ are sampled from the $K$ groups with probabilities $p_{1},\dots,p_{K}$, as in (4), then $\widehat{p}_{k}$ is an estimate of $p_{k}$ for each $k$. Then, for any $X$ with $X^{0}=k$, the weight for the covariate shift under this model is given by $w(X)=q_{k}/p_{k}$; we can estimate this weight with $\widehat{w}(X)=q_{k}/\widehat{p}_{k}$. In this comparison, we will assume for simplicity that $n_{k}>0$ for all $k$, i.e., all groups are observed in the training set—otherwise the function $\widehat{w}$ would not be well-defined. By definition, we have $\sum_{i\in[n]}\widehat{w}(X_{i})=n$. Then rewriting the definition of $\widehat{q}$ one more time,
@@
As we will see below, the prediction interval defined here in (6) leads to a coverage guarantee that is slightly smaller than the target level $1-\alpha$. Interestingly, however, this cannot be fixed by returning to the slightly more conservative WCP method given in (2). We choose to work with this present formulation, (6), since avoiding the additional weight on $\delta_{+\infty}$ leads to a simpler construction and a (slightly) less conservative method without any additional loss of coverage in the theoretical guarantee. Of course, any coverage guarantee that can be established for the prediction interval defined in (6) will also hold for the version of WCP that adds a weight to $\delta_{+\infty}$, since this leads to a strictly more conservative prediction interval.
@@ 3.2 Estimation error in the group-wise setting: applying prior results
However, even in this setting of a finite-dimensional unknown parameter, existing coverage guarantees provide a fairly pessimistic view and imply that the loss of coverage may be severe. In particular, if our estimate of the probabilities $p_{1},\dots,p_{K}$ (which determine the relative proportions of the $K$ groups in the training distribution $P$) is based on a sample of size $\mathcal{O}(n)$, then we would expect that each $p_{k}$ is estimated with error scaling as $\mathcal{O}(\sqrt{p_{k}/n})$ (due to the variance of a Binomial distribution); equivalently, the relative error in estimating $p_{k}$ scales as $\mathcal{O}(1/\sqrt{np_{k}})$. For example, if $p_{k}\equiv 1/K$ (i.e., the true distribution places equal weight on each group), then the relative error in estimating each $p_{k}$ scales as $\mathcal{O}(\sqrt{K/n})$, and so the coverage loss term in (Lei and Candès, 2021)’s guarantee (3) can be expected to scale as
@@
To demonstrate this intuition more concretely, we compute (Lei and Candès, 2021)’s coverage guarantee (3) empirically, across a range of sample sizes $n=100,110,120,\dots,1000$, in three regimes: a constant number of groups, $K=10$; a slowly increasing number of groups, $K=\lfloor\sqrt{n}\rfloor$; and a proportional number of groups, $K=n/10$. In each case, the probabilities are given by $p_{k}=q_{k}\equiv 1/K$ for all groups $k\in[K]$, for both the training and test distributions. The estimated weight function $\widehat{w}$ is then given by $\widehat{w}(X)=q_{k}/\widehat{p}_{k}$ for any data point with $X^{0}=k$, where $q_{k}\equiv 1/K$ is known, while $\widehat{p}_{k}$ is given by the empirical proportion of group $k$ in a pretraining sample of size $n$. To avoid dividing by zero, we add a +1 adjustment to each group’s count—that is, $\widehat{p}_{k}=\frac{n_{k}+1}{n+K}$, where $n_{k}$ is the number of samples observed in group $k$ within the training set of size $n$ (since the true proportions $p_{k}$ are uniform, this adjustment can only make the estimates more accurate).
@@
Figure 1: (Lei and Candès, 2021)’s WCP coverage guarantee (3) for varying numbers of groups. The horizontal line marks the target coverage level 90% (i.e., $\alpha=0.1$). See Section 3.2 for details.